What are the solutions to this equation (2x+3)^2=10
Answer by TimothyLamb(4379)

---
(2x+3)^2=10
---
(2x+3)(2x+3) = 10
4xx + 12x + 9 = 10
4xx + 12x - 1 = 0
---
the above quadratic equation is in standard form, with a=4, b=12, and c=-1
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
4 12 -1
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
x = 0.0811388301
x = -3.08113883
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php